# Assessment of Spatio-Temporal Predictors in the Presence of Missing and Heterogeneous Data

### v2

@@ 4 AZ-whiteness test
The analysis presented in this paper leverages the AZ-whiteness test (Zambon and Alippi, 2022), a statistical test designed to detect the presence of both autocorrelation and cross-correlation. The test is defined over hypotheses
$\begin{cases}H_{0}:&\text{All pairs }(\boldsymbol{\mathbf{r}}_{t,v},\boldsymbol{\mathbf{r}}_{\tau,u})\text{ are uncorrelated,}\\ H_{1}:&\text{At least a pair }(\boldsymbol{\mathbf{r}}_{t,v},\boldsymbol{\mathbf{r}}_{\tau,u})\text{ is correlated,}\end{cases}$ (11)
for $(t,v)\neq(\tau,u)$, and rejects the null hypothesis $H_{0}$ whenever the absolute value of statistic $C_{\lambda}(\boldsymbol{\mathbf{g}}^{*})$ exceeds a threshold $\gamma$ predetermined according to a user-defined significance level $\alpha$
@@
The AZ-whiteness test statistic $C_{\lambda}(\boldsymbol{\mathbf{g}}^{*})\in\mathbb{R}$ is computed on the prediction residuals $\{\boldsymbol{\mathbf{r}}_{t,v}:\forall\ t,v\}$ with reference to the spatio-temporal multiplex graph $\boldsymbol{\mathbf{g}}^{*}$. More in detail, the test statistic is defined as
@@
respectively. Scalar quantities $\widetilde{C}_{\rm sp}$ and $\widetilde{C}_{\rm tm}$ are weighted sums assessing the spatial and temporal correlation, respectively, by accounting for positive and negative signs
@@
between residual vectors $\boldsymbol{\mathbf{r}}_{t,v}$ and $\boldsymbol{\mathbf{r}}_{\tau,u}$; sign function ${\rm sgn}(a)$ equals $-1,0$ or $1$ depending on whether $a$ is negative, null, or positive. Scalar $w_{\rm tm}$ is a single weight assigned to all temporal edges and set such that the normalization term $W_{\rm tm}\doteq|E_{\rm tm}|\cdot w_{\rm tm}^{2}$ equals $W_{\rm sp}$
@@
Figure 2: A section view of the spatio-temporal graph $\boldsymbol{\mathbf{g}}^{*}$ from Sec. 3.2. Each node is associated with a residual vector $\boldsymbol{\mathbf{r}}_{t,v}$ and each edge – either spatial (red) or temporal (blue) – is associated with a sign (16).
The distinction between spatial and temporal contributions is also highlighted in Figure 2. The intuition behind the test statistic (13) is that large positive values of $C_{\lambda}(\boldsymbol{\mathbf{g}}^{*})$ indicate the presence of adjacent residuals similarly oriented, thus suggesting correlation among variables; see also Figure 3 discussed below. Similarly, $C_{\lambda}(\boldsymbol{\mathbf{g}}^{*})\ll 0$ provides evidence of negative correlation.
Identifying a value of $\gamma$ for any choice of $\alpha$ is possible thanks to Theorem 1 showing that, asymptotically with the number of edges, $C_{\lambda}(\boldsymbol{\mathbf{g}}^{*})$ follows a standard Gaussian distribution irrespective of the distribution of the residuals and without requesting them to be identically distributed.
@@
    $\mathbb{E}_{\boldsymbol{\mathbf{r}}_{t,v}}\left[{\rm sgn}\left(\bar{\boldsymbol{\mathbf{r}}}^{\top}\boldsymbol{\mathbf{r}}_{t,v}\right)\right]=\boldsymbol{\mathbf{0}}$ for all $\bar{\boldsymbol{\mathbf{r}}}\in\mathbb{R}^{d_{y}}\setminus\{\boldsymbol{\mathbf{0}}\}$, and $v_{t}\in V^{*}$,
@@
The assumptions of Theorem 1 are very mild, thanks also to the choice of assessing correlation by means of signs (16). More in detail, A1 is requested to meet the null hypothesis, while A2 and A3 are straightforwardly resolved in most scenarios, if not already fulfilled (Zambon and Alippi, 2022). In particular, for scalar residuals, A2 reduces to requesting that the residual median is null. Lastly, A3 requests positive bounded weights, as they are assumed here to encode the strength of the dependency between the related nodes. Whenever no weights are given, or they do not fulfill A3, we recommend setting all weights to $1$.
@@ 5 AZ-analysis of residuals
The AZ-analysis introduced in this paper addresses questions Q1Q3 by inspecting families $\{\boldsymbol{\mathbf{s}}_{i}\}$ of subsets $\boldsymbol{\mathbf{s}}_{i}$ of all possible pairs of residuals in $\boldsymbol{\mathbf{r}}$ and identifying those that display more prominent evidence of correlation according to scoring function $c_{\lambda}(\cdot)$, related to $C_{\lambda}(\cdot)$ as explained in next subsection. It is convenient to refer subsets $\boldsymbol{\mathbf{s}}_{i}$ as subgraphs of the fully connected graph connecting all pairs of observations in $\boldsymbol{\mathbf{r}}$. Given suitably designed subgraphs, we compute and compare correlation scores $c_{\lambda}(\boldsymbol{\mathbf{s}}_{i})$ for different subgraphs $\boldsymbol{\mathbf{s}}_{i}$, having in mind that $|c_{\lambda}(\boldsymbol{\mathbf{s}}_{i})|>|c_{\lambda}(\boldsymbol{\mathbf{s}}_{j})|$ implies the existence of a larger residual correlation in $\boldsymbol{\mathbf{s}}_{i}$ than in $\boldsymbol{\mathbf{s}}_{j}$. Although subgraphs $\{\boldsymbol{\mathbf{s}}_{i}\}$ may differ in size, we show that the associated scores $\{c_{\lambda}(\boldsymbol{\mathbf{s}}_{i})\}$ can be compared and ranked.
Figure 3: The figure compares the value of of statistics $C_{\lambda}(\cdot)$ defined in (13) (left) with that of the correlation scores $c_{\lambda}(\cdot)$ of (20) (right). Different levels of residual correlation and number $|E|$ of graph edges are considered. Each color corresponds to a different number of edges. Solid lines represent the expected value of the score estimated over 100 repeated experiments, dashed lines the standard deviation, shaded area the interquartile interval.
@@ 5.1 Correlation scores
Under the null hypothesis, test statistic $C_{\lambda}(\boldsymbol{\mathbf{s}})$ asymptotically follows a standard Gaussian distribution, regardless of the size of the graph $\boldsymbol{\mathbf{s}}$; see Theorem 1. This implies that all sample means should be centered around zero, except for those indicating evidence of correlation, while maintaining the same standard deviation, thereby ensuring a certain degree of comparability.
@@
The left panel of Figure 3 illustrates values of $C_{\lambda}(\boldsymbol{\mathbf{s}})$ as function of the number of edges $|E_{\boldsymbol{\mathbf{s}}}|$ in $\boldsymbol{\mathbf{s}}$ and the Pearson correlation between residuals; detailed description of the experiment is provided in Appendix A. The figure illustrates that when no correlation exists among residuals, all scores are centered around zero with a standard deviation of 1, regardless of the number of edges considered. As correlation increases, each curve – corresponding to a different number of edges – edges—exhibits a monotonic increasing trend. However, we also observe that scores based on a larger number of edges are more effective in identifying correlated residuals. This suggests introducing a statistic that enables the comparison of score values independent of the number of edges used in their computation.
The maximum value of each curve is $\sqrt{|E_{\boldsymbol{\mathbf{s}}}|}$. In fact, when all signs in (13) equal 1, the test statistic [^1] For brevity, here $\boldsymbol{\mathbf{s}}$ is a generic graph with no explicit distinction between spatial and temporal edges. Accordingly, $\widetilde{C}$ is the weighted sum across all edges in set $E_{\boldsymbol{\mathbf{s}}}$ (regardless of whether they are spatial or temporal), $\lambda$ is incorporated in the weights $\boldsymbol{\mathbf{w}}$ and the denominator of (13) reduces to $\lVert\boldsymbol{\mathbf{w}}\rVert_{2}$. is
@@
$C_{\lambda}(\boldsymbol{\mathbf{s}})=\frac{\widetilde{C}}{\sqrt{\lVert\boldsymbol{\mathbf{w}}\rVert_{2}^{2}}}=\frac{\lVert\boldsymbol{\mathbf{w}}\rVert_{1}}{\lVert\boldsymbol{\mathbf{w}}\rVert_{2}};$ (19)
note that in the experiment yielding to Figure 3 all edge weights are equal to 1, therefore $\lVert\boldsymbol{\mathbf{w}}\rVert_{1}=|E_{\boldsymbol{\mathbf{s}}}|$ and $\lVert\boldsymbol{\mathbf{w}}\rVert_{2}=\sqrt{|E_{\boldsymbol{\mathbf{s}}}|}$. Following this observation, we scale statistic $C_{\lambda}(\cdot)$ to yield values in the $[-1,1]$ interval as
@@
$c_{\lambda}(\boldsymbol{\mathbf{s}})\doteq C_{\lambda}(\boldsymbol{\mathbf{s}})\frac{\lVert\boldsymbol{\mathbf{w}}\rVert_{2}}{\lVert\boldsymbol{\mathbf{w}}\rVert_{1}}=\frac{\widetilde{C}}{\lVert\boldsymbol{\mathbf{w}}\rVert_{1}},$ (20)
and name $c_{\lambda}(\cdot)$ correlation score.
@@
The right panel of Figure 3 shows the behavior of scores $c_{\lambda}(\cdot)$. While they share the same advantages as $C_{\lambda}(\cdot)$, where magnitude increases as correlation increases, we also observe that their expected value remains the same regardless of the number of edges. Accordingly, we suggest to inspect those subgraphs $\boldsymbol{\mathbf{s}}_{i}$ associated with $|C_{\lambda}(\boldsymbol{\mathbf{s}}_{i})|\gg 0$ (evidencing correlation), starting from those with large $|c_{\lambda}(\boldsymbol{\mathbf{s}}_{i})|$ in decreasing order.
@@ 5.2 Subgraphs as sets of edges
Note that the number of correlated pairs of residuals is, in principle, quadratic in the number of residuals $\boldsymbol{\mathbf{r}}$ in (5), $N\,T$, where $T$ is the length of the time series and $N$ the average number of available sensors. More precisely, there are $N\,T$ observed residuals – corresponding to the nodes $v_{t}\in V^{*}$ of $\boldsymbol{\mathbf{g}}^{*}$ – which leads to $(N\,T)(N\,T-1)/2$ (possibly) correlated node pairs. Indeed, considering all possible subsets of node pairs is unfeasible, as their number is about $2^{{N^{2}\,T^{2}/2}}$. We deal with this complexity problem in two steps.
First, we confine the analysis to node pairs corresponding to edges in $\boldsymbol{\mathbf{g}}^{*}$, which amounts to considering all possible subgraphs of $\boldsymbol{\mathbf{g}}^{*}$. As the relational information encoded in multiplex $\boldsymbol{\mathbf{g}}^{*}$ is relevant to solving the downstream task, the node pairs are those that are more likely to exhibit correlation in the associated residuals. The inspection reduces to $2^{E_{*}}$ subgraphs, i.e., approximately $2^{N\,\delta\,T/2}$, with $\delta$ the average node degree in the graph sequence incremented by 2 to account for temporal edges. Second, as computing and inspecting scores for all $2^{E_{*}}$ subsets is still impractical, we propose principled families of subgraphs designed to address the three key questions Q1, Q2, and Q3 in Sections 5.3 to 5.6.
Figure 4: Top) Generic subgraphs $\boldsymbol{\mathbf{s}}_{v},\boldsymbol{\mathbf{s}}_{u}$ involved in the computation of node scores $c_{\lambda}(v)$ and $c_{\lambda}(u)$. Center) Generic subgraphs $\boldsymbol{\mathbf{s}}_{t},\boldsymbol{\mathbf{s}}_{\tau}$ involved in the computation of temporal scores $c_{\lambda}(t)$ and $c_{\lambda}(\tau)$. Bottom) Generic subgraphs $\boldsymbol{\mathbf{s}}_{t,v},\boldsymbol{\mathbf{s}}_{\tau,u}$ involved in the computation of temporal scores $c_{\lambda}(v,t)$ and $c_{\lambda}(\tau,u)$.

### v3

@@ 4 AZ-whiteness test
The analysis presented in this paper leverages the AZ-whiteness test (Zambon and Alippi, 2022), a statistical test designed to detect the presence of both autocorrelation and cross-correlation in data streams. The test is defined over hypotheses
[…]
for $(t,v)\neq(\tau,u)$, and rejects the null hypothesis $H_{0}$ whenever the absolute value of the statistic $C_{\lambda}(\boldsymbol{\mathbf{g}}^{*})$ exceeds a threshold $\gamma$. This threshold is predetermined according to a user-defined significance level $\alpha$
@@
The AZ-whiteness test statistic $C_{\lambda}(\boldsymbol{\mathbf{g}}^{*})\in\mathbb{R}$ is computed on the prediction residuals $\{\boldsymbol{\mathbf{r}}_{t,v}:\forall\ t,v\}$ with reference to the spatio-temporal multiplex graph $\boldsymbol{\mathbf{g}}^{*}$. In more detail, the test statistic is defined as
@@
respectively. Scalar quantities $\widetilde{C}_{\rm sp}$ and $\widetilde{C}_{\rm tm}$ are weighted sums assessing spatial and temporal correlation, respectively, by accounting for positive and negative signs
@@
between residual vectors $\boldsymbol{\mathbf{r}}_{t,v}$ and $\boldsymbol{\mathbf{r}}_{\tau,u}$; sign function ${\rm sgn}(a)$ equals 1,0-1,0 or $1$ depending on whether $a$ is negative, null, or positive. Scalar $w_{\rm tm}$ is a single weight assigned to all temporal edges and set such that the normalization term $W_{\rm tm}\doteq|E_{\rm tm}|\cdot w_{\rm tm}^{2}$ equals $W_{\rm sp}$
@@
Figure 2: A section view of the spatio-temporal graph $\boldsymbol{\mathbf{g}}^{*}$ from Section 3.2. Each node is associated with a residual vector $\boldsymbol{\mathbf{r}}_{t,v}$ and each edge – either spatial (red) or temporal (blue) – is associated with a sign (16).
The distinction between spatial and temporal contributions is also highlighted in Figure 2. The intuition behind the test statistic (13) is that large positive values of $C_{\lambda}(\boldsymbol{\mathbf{g}}^{*})$ indicate similarly oriented adjacent residuals, suggesting positive correlation among variables (see also Figure 3). Likewise, strongly negative values of $C_{\lambda}(\boldsymbol{\mathbf{g}}^{*})\ll 0$ indicate negative correlation.
Identifying a value of $\gamma$ for any choice of $\alpha$ is possible thanks to Theorem 1. This theorem shows that, asymptotically with the number of edges, $C_{\lambda}(\boldsymbol{\mathbf{g}}^{*})$ follows a standard Gaussian distribution. Importantly, this result holds irrespective of the distribution of the residuals and does not require them to be identically distributed.
@@
    $\mathbb{E}_{\boldsymbol{\mathbf{r}}_{t,v}}\left[{\rm sgn}\left(\bar{\boldsymbol{\mathbf{r}}}^{\top}\boldsymbol{\mathbf{r}}_{t,v}\right)\right]=0$ for all $\bar{\boldsymbol{\mathbf{r}}}\in\mathbb{R}^{d_{y}}\setminus\{\boldsymbol{\mathbf{0}}\}$ and $v_{t}\in V^{*}$,
@@
The assumptions of Theorem 1 are very mild. This is partly due to the choice of assessing correlation using signs, as defined in (16). More specifically, assumption A1 is required to satisfy the null hypothesis. In contrast, A2 and A3 are straightforwardly resolved in most scenarios, if not already fulfilled (Zambon and Alippi, 2022). In particular, for scalar residuals, A2 reduces to requesting that the residual median is zero. Lastly, A3 requests weights to be positive and bounded, as they are assumed to encode the strength of the dependency between nodes. Whenever no weights are given, or if they do not satisfy A3, we suggest considering setting all weights to $1$ or monotonically transforming them.
@@ 5 AZ-analysis of residuals
The AZ-analysis introduced in this paper addresses questions Q1Q3 by inspecting families $\{\boldsymbol{\mathbf{s}}_{i}\}$ of subsets of all possible pairs of residuals in $\boldsymbol{\mathbf{r}}$. The goal is to identify those subsets that display stronger evidence of correlation, as measured by a scoring function $c_{\lambda}(\cdot)$, which is related to $C_{\lambda}(\cdot)$ as explained in the next subsection. It is convenient to refer subsets $\boldsymbol{\mathbf{s}}_{i}$ as subgraphs of the fully connected graph over all observations in $\boldsymbol{\mathbf{r}}$. Given suitably designed subgraphs, we compute and compare correlation scores $c_{\lambda}(\boldsymbol{\mathbf{s}}_{i})$ for different subgraphs $\boldsymbol{\mathbf{s}}_{i}$, having in mind that $|c_{\lambda}(\boldsymbol{\mathbf{s}}_{i})|>|c_{\lambda}(\boldsymbol{\mathbf{s}}_{j})|$ implies the existence of a larger residual correlation in $\boldsymbol{\mathbf{s}}_{i}$ than in $\boldsymbol{\mathbf{s}}_{j}$. Although subgraphs $\{\boldsymbol{\mathbf{s}}_{i}\}$ may differ in size, we show that the associated scores $\{c_{\lambda}(\boldsymbol{\mathbf{s}}_{i})\}$ can be compared and ranked.
The remainder of this section is structured as follows. We introduce the notion of correlation scores and explain how they enable comparisons across different residual subsets in Section 5.1. We then discuss the practical challenge of handling the large number of possible subgraphs and describe strategies for defining meaningful subgraph families in Section 5.2. Subsequent subsections (Sections 5.35.6) address the three key questions Q1Q3 related to the identification of residual correlations. Finally, we offer practical considerations and implementation guidance in Section 5.7.
Figure 3: The figure compares the value of statistics $C_{\lambda}(\cdot)$ defined in (13) (left panel, 3(a)) with that of the correlation scores $c_{\lambda}(\cdot)$ of (20) (right panel, 3(b)). Different levels of residual correlation and the number $|E|$ of graph edges are considered. Each color corresponds to a different number of edges. Solid lines represent the expected value of the score estimated over 100 repeated experiments, dashed lines the standard deviation, shaded area the interquartile interval.
@@ 5.1 Correlation scores
Under the null hypothesis, test statistic $C_{\lambda}(\boldsymbol{\mathbf{s}})$ asymptotically follows a standard Gaussian distribution, regardless of the size of the graph $\boldsymbol{\mathbf{s}}$; see Theorem 1. This implies that all sample means should be centered around zero, except for those indicating evidence of correlation. Additionally, the test maintains the same standard deviation, ensuring a certain degree of comparability.
@@
The left panel of Figure 3 illustrates values of $C_{\lambda}(\boldsymbol{\mathbf{s}})$ as a function of the total number of edges $|E_{\boldsymbol{\mathbf{s}}}|$ in $\boldsymbol{\mathbf{s}}$ and the Pearson correlation between residuals. A detailed description of the experiment is provided in A. The figure illustrates that when no correlation exists among residuals, all scores are centered around zero with a standard deviation of 1, regardless of the number of edges considered. As correlation increases, each curve – corresponding to a different number of edges – exhibits a monotonic increasing trend. However, we also observe that scores based on a larger number of edges are more effective in identifying correlated residuals. This suggests introducing a statistic that enables the comparison of score values independent of the number of edges used in their computation.
The maximum value of each curve in Figure 3 is $\sqrt{|E_{\boldsymbol{\mathbf{s}}}|}$. In fact, in the setting of the experiment, all weights (accounting for the factors $\lambda$ and $1-\lambda$ too) and all signs in (13) are equal to 1. Therefore, the test statistic simplifies to
$C_{\lambda}(\boldsymbol{\mathbf{s}})=\frac{\sum_{e\in E_{\rm sp}}1+\sum_{e\in E_{\rm tm}}1}{\sqrt{\sum_{e\in E_{\rm sp}}1^{2}+\sum_{e\in E_{\rm tm}}1^{2}}}=\frac{|E_{\boldsymbol{\mathbf{s}}}|}{\sqrt{|E_{\boldsymbol{\mathbf{s}}}|}}.$ (19)
More generally, the following lemma holds.
Lemma 2.
Denote $\boldsymbol{\mathbf{w}}_{\boldsymbol{\mathbf{s}},\lambda}=\{\lambda w_{e}:e\in E_{{\rm sp}}\}\cup\{(1-\lambda)w_{e}:e\in E_{{\rm tm}}\}$ the set of all spatial and temporal weights of $\boldsymbol{\mathbf{s}}$, adjusted according to $\lambda$. The maximum and minimum values of $C_{\lambda}(\boldsymbol{\mathbf{s}})$ are $\pm\lVert\boldsymbol{\mathbf{w}}_{\boldsymbol{\mathbf{s}},\lambda}\rVert_{1}/\lVert\boldsymbol{\mathbf{w}}_{\boldsymbol{\mathbf{s}},\lambda}\rVert_{2}$.
Proof.
The numerator and denominator in (13) satisfy
@@
$\left|\lambda\widetilde{C}_{\rm sp}+(1-\lambda)\widetilde{C}_{\rm tm}\right|$ $\leq\sum_{e\in E_{{\rm sp}}}|\lambda w_{e}|+\sum_{e\in E_{{\rm tm}}}|(1-\lambda)w_{e}|=\lVert\boldsymbol{\mathbf{w}}_{\boldsymbol{\mathbf{s}},\lambda}\rVert_{1}$
$\lambda^{2}W_{{\rm sp}}+(1-\lambda)^{2}W_{{\rm tm}}$ $=\lVert\boldsymbol{\mathbf{w}}_{\boldsymbol{\mathbf{s}},\lambda}\rVert_{2}^{2}.$
In particular, the bounds are tight and are achieved when the signs are either all positive (+1+1) or all negative (1-1). Therefore, $|C_{\lambda}(\boldsymbol{\mathbf{s}})|\leq{\lVert\boldsymbol{\mathbf{w}}_{\boldsymbol{\mathbf{s}},\lambda}\rVert_{1}}/{\lVert\boldsymbol{\mathbf{w}}_{\boldsymbol{\mathbf{s}},\lambda}\rVert_{2}}$. ∎
Based on Lemma 2 above, we scale the statistic $C_{\lambda}(\cdot)$ so that its values fall within the interval [1,1][-1,1] as
@@
$c_{\lambda}(\boldsymbol{\mathbf{s}})\doteq C_{\lambda}(\boldsymbol{\mathbf{s}})\frac{\lVert\boldsymbol{\mathbf{w}}_{\boldsymbol{\mathbf{s}},\lambda}\rVert_{2}}{\lVert\boldsymbol{\mathbf{w}}_{\boldsymbol{\mathbf{s}},\lambda}\rVert_{1}}=\frac{\lambda\widetilde{C}_{\rm sp}+(1-\lambda)\widetilde{C}_{\rm tm}}{\lambda\sum_{e\in E_{{\rm sp}}}w_{e}+(1-\lambda)\sum_{e\in E_{{\rm tm}}}w_{e}};$ (20)
call $c_{\lambda}(\cdot)$ the correlation scores.
@@
The right panel of Figure 3 shows the behavior of scores $c_{\lambda}(\cdot)$. While scores share the same advantages as $C_{\lambda}(\cdot)$, in that their magnitude increases with the residual correlation, we also observe that their expected values do not depend on the number of edges. In particular, $\mathbb{E}[c_{\lambda}(\boldsymbol{\mathbf{s}})]=0$ under Assumptions A1 and A2. Accordingly, we suggest inspecting those subgraphs $\boldsymbol{\mathbf{s}}_{i}$ associated with $|C_{\lambda}(\boldsymbol{\mathbf{s}}_{i})|\gg 0$ (evidencing correlation, as discussed in Section 4), starting from those with large $|c_{\lambda}(\boldsymbol{\mathbf{s}}_{i})|$ in decreasing order; indeed, checking the presence of correlation first avoids inspecting and comparing scores whose values do not reflect any significant correlation.
@@ 5.2 Subgraphs as sets of edges
Note that the number of potentially correlated pairs of residuals is, in principle, quadratic in the number of residuals $\boldsymbol{\mathbf{r}}_{t,v}$ in (5). Denoting by $T$ the length of the time series and by $N$ the average number of available sensors, there are NTN\,T observed residuals, which leads to (NT)(NT1)/2(N\,T)(N\,T-1)/2 pairs. Indeed, considering all possible subsets of node pairs is, in practice, infeasible, since their number is approximately $2^{{N^{2}\,T^{2}/2}}$. We address this complexity problem in two steps.
First, we confine the analysis to node pairs corresponding to edges in $\boldsymbol{\mathbf{g}}^{*}$, which amounts to considering all possible subgraphs of $\boldsymbol{\mathbf{g}}^{*}$. The relational information encoded in multiplex $\boldsymbol{\mathbf{g}}^{*}$ is relevant for solving the downstream task. Therefore, the node pairs represented by its edges are more likely to exhibit correlation in the associated residuals. The inspection reduces to $2^{|E_{*}|}$ subgraphs, i.e., approximately $2^{N\,\delta\,T/2}$, with $\delta$ the average node degree in the graph sequence incremented by 2 to account for temporal edges. Second, even analyzing all $2^{E_{*}}$ subsets remains impractical. Therefore, we propose principled families of subgraphs, each designed to address the three key questions Q1, Q2, and Q3 in Sections 5.35.6.
Figure 4: Top right) Generic subgraphs $\boldsymbol{\mathbf{s}}_{t},\boldsymbol{\mathbf{s}}_{\tau}$ involved in the computation of time scores $c_{\lambda}(t)$ and $c_{\lambda}(\tau)$. Center right) Generic subgraphs $\boldsymbol{\mathbf{s}}_{v},\boldsymbol{\mathbf{s}}_{u}$ involved in the computation of node scores $c_{\lambda}(v)$ and $c_{\lambda}(u)$. Bottom right) Generic subgraphs $\boldsymbol{\mathbf{s}}_{t,v},\boldsymbol{\mathbf{s}}_{\tau,u}$ involved in the computation of local spatio-temporal scores $c_{\lambda}(v,t)$ and $c_{\lambda}(\tau,u)$. Left) Convenient arrangement of the scores for visualization purposes; see subsequent experimental sections.